Chair: Simin

Next meeting chair: Katie

Attended: Jamie, Jasper, Earl, Mas, Prad, Lim

Apologies:

General Issues

  • Yg: We have bad pointing offset, a temporary fix “azeloff=0d,0.13443d” is in place, please do not change/reset this. This fix will need to be applied to the follow-on T2x experiment during the setup.

It is fixed for now. We need to look at the results of pointing-check during the experiment setups and let Jamie know about large pointing offsets above 0.03 or 0.04 degrees.

Handover Issues

AOV059

Hb

  • FS unresponsive, antenna coms problems. rebooting pcfshb (JMc). fs fell over again! Antenna comms bad, gpib not talking etc (warren).
  • fs is upgraded and some changes happened like the “antenna=open” is working when “open” is in uppercase –> “antenna=OPEN”

Ke

  • Sampler 3 (X-band, X-pol) reads as bad. Stopping to reconfigure (JMc).
  • connection problem again. “antenna=open” & “antenna=operate” for several times and finally reset fs (Simin).